# Can Diffusion Models Provide Rigorous Uncertainty Quantification for Bayesian Inverse Problems?

### v1

@@ 5.3 STYLIZED PHASE RETRIEVAL STUDY
Table 4 shows the average errors in estimation of the mean and variance, as well as average CMD and MMD errors, for all nine BIPSDA methods tested. Here we first note that, when the score is known analytically, the TC based approaches perform better than the TU and ODE based approaches, and on some trials produce strong performance, as evidenced by the provided interdecile ranges. This indicates that the TC approaches have some promise if a reliable approximation of the second-order score can be estimated using, e.g. the methods in [30]. However, there is wide variance in performance across the trials, and in general the TC approaches cannot fully overcome the issues observed with the TU approaches in Figure 2. All of the BIPDSA algorithms we tested thus face performance issues in this problem context.
@@ 6 DISCUSSION AND CONCLUSION
Diffusion annealing based inverse problem solvers is an emerging class of algorithms for solving Bayesian inverse problems using a diffusion model trained on the prior distribution. In this work, we introduced a general framework, Bayesian Inverse Problem Solvers through Diffusion Annealing (BIPSDA), for this class of algorithms. This framework has two key design choices—the denoising distribution approximation and the sampler for the prediction distribution—that can be “mixed and matched” in a flexible manner. The large algorithmic design space of BIPSDA includes the previously introduced DAPS [16] and DiffPIR [17] algorithms, which have both achieved state-of-the-art performance on different image reconstruction problems.
We analyzed the performance of nine algorithms within the framework, including the DAPS algorithm, DiffPIR algorithm, and seven novel algorithms, on a suite of model problems where the posterior density is known. Here the novel algorithms were realized through combinations of ideas from DAPS and DiffPIR. We also incorporated the Tweedie Correlated (TC) technique [30] for modeling the denoising distribution and a randomize-then-optimize (RTO) technique [18, 37, 19] for fast, approximate sampling from the prediction distribution. The TC technique was previously introduced in the context of the hijacking class of diffusion-based Bayesian inverse problem solvers [26], but has not been employed in the diffusion annealing context. The RTO-based idea, to the best of the author’s knowledge, has not previously been employed in the context of diffusion modeling.
@@ 6.1 RESULTS ANALYSIS
The results provide insights into both the impact of specific design choices with our framework and the performance of diffusion model based inverse problem solvers generally. Regarding the choice of denoising distribution approximation, we note that approaches that use the TC technique provide the best performance on the two non-linear inverse problems we tested when the prior score is assumed to be known analytically. However, in our implementation this technique is currently only applicable when the score is known exactly. The training of an auxiliary neural network model to approximate higher-order moments of the data distribution is required in the learned score setting [30]. It is also worth noting that in general, the TU technique, which uses an unbiased estimator of the mean of the denoising distribution, performs better than the ODE technique, which uses a biased estimate.
Regarding the sampler for the prediction distribution, the results demonstrate that the MAP estimation based approaches, despite lacking firm theoretical foundations, perform well in recovering the global structure of the posterior distribution. However, they systematically underestimate the variance of each of the posterior modes, which is unsurprising given that the MAP approaches do not actually sample from the prediction distribution. The RTO technique, which is equivalent to exact sampling from the prediction distribution when the likelihood function is linear-Gaussian [38], partially resolves this issue and provides better performance than the MAP based approaches on the stylized x-ray tomography problem and inpainting problems. Finally, the Langevin dynamics based approaches work well when the posterior is unimodal (low noise regime of the stylized impainting problem and stylized x-ray tomography problem). However, in the case of multimodal posteriors (as in the high noise regime of the stylized inpainting problem), Langevin dynamics struggles to properly incorporate the measurement information. Further, on the phase retrieval problem, Langevin dynamics without Metropolis correction completely fails, and the incorporation of a Metropolis adjustment and preconditioning was required to obtain competitive performance.
@@ 6.2 LIMITATIONS AND FUTURE DIRECTIONS
First, it is of interest to develop an extension of TC technique to the case in which the score is not known analytically but is learned from data. This could potentially be accomplished through regularization of the TC estimator and the use of an additional neural network trained to match the second-order moments of the data distribution [30]. Second, the algorithms in our framework were tested on stylized model problems that enable principled performance analysis but have low dimensionality. While previously established algorithms within our framework have been shown to produce high quality samples for large-scale problems (including image reconstruction), it is of interest to systematically evaluate the performance BIPSDA framework in these high-dimensional settings.
It is also of interest to improve the performance of BIPSDA algorithms on challenging non-linear problems like phase retrieval. Here we first note that in this study there were cases where posterior samples numerically diverged due to large error in the learned score in low probability regions of the prior. This could potentially be addressed through the incorporation of second-order score model derivative information into the score training, which has been shown to improve model accuracy in low-density regions in other problem contexts [47]. Further, even with the score known analytically, the BIPSDA algorithms performed poorly on some of the trials, and additional algorithmic innovations are required to address this. In particular, it is of interest to explore modifications to the proposed RTO technique, such as apodizing the RTO noise perturbation in early BIPDSA iterations or metropolizing the RTO sampling, to improve performance on challenging problems like this one.
@@
7 MEAN ESTIMATION WITH PROBABILITY FLOW ODES
In this appendix, we analyze the relationship between the Tweedie’s formula and probabilify flow ODE based approaches for estimation the mean of the denoising distribution $\mathbb{E}_{0\mid t}[\mathbf{m}(0)\mid\mathbf{m}(t)]$.
In the Tweedie’s formula based approach, the estimated of the predicted mean, denoted $\mathbf{m}_{\mathrm{tweedie}}$, is computed using the pretrained score model as
$\mathbf{m}_{\mathrm{tweedie}}=\mathbf{m}(t)+[\sigma^{2}(t)-\sigma^{2}(0)]s_{\boldsymbol{\theta}^{*}}(\mathbf{m}(t),t).$ (11)
Since the conditional distribution of $\mathbf{m}(t)$ given $\mathbf{m}(0)$ is Gaussian, Tweedie’s formula is exact and
$\mathbb{E}_{0\mid t}[\mathbf{m}(0)\mid\mathbf{m}(t)]$ $=\mathbf{m}(t)$
$\quad+[\sigma^{2}(t)-\sigma^{2}(0)]\nabla_{\mathbf{m}(t)}\log\pi_{t}(\mathbf{m}(t)).$
@@ 7 MEAN ESTIMATION WITH PROBABILITY FLOW ODES
As can be seen, the ODE and Tweedie’s formula based approaches both update $\mathbf{m}(t)$ in the direction of $s_{\boldsymbol{\theta}^{*}}(\mathbf{m}(t),t)$. However, the magnitude of the update differs by a factor of $d=t\sigma(t)\dot{\sigma}(t)/(\sigma^{2}(t)-\sigma^{2}(0))$, which in general will not be equal to one, and so therefore $\mathbf{m}_{\mathrm{ode}}\neq\mathbf{m}_{\mathrm{tweedie}}$. However, under the particular choice of parameterization $\sigma(t)=t$, then $d\approx 1$ and the updates approximately coincide.
Let us now consider the same diffusion model parameterization, i.e., $\sigma(t)=t$, but with two steps used in the reverse Euler discretization. Then we have that
$\mathbf{m}_{\mathrm{ode}}=\mathbf{m}(t_{m})+t_{m}^{2}s_{\boldsymbol{\theta}^{*}}(\mathbf{m}(t_{m}),t_{m}),$
with
$\mathbf{m}(t_{m})=\mathbf{m}(t)+t(t-t_{m})s_{\boldsymbol{\theta}^{*}}(\mathbf{m}(t_{m}),t_{m}),$
where $t_{m}\in(0,t)$ is the discretization point. Combining the two equations above, we have that
$\mathbf{m}_{\mathrm{ode}}$ $=\mathbf{m}(t)+t(t-t_{m})s_{\boldsymbol{\theta}^{*}}(\mathbf{m}(t_{m}),t_{m})$
$\quad+t_{m}^{2}s_{\boldsymbol{\theta}^{*}}(\mathbf{m}(t)+t(t-t_{m})s_{\boldsymbol{\theta}^{*}}(\mathbf{m}(t_{m}),t_{m}),t_{m}),$
and it is clear that in this setting, $\mathbf{m}_{\mathrm{ode}}\neq\mathbf{m}_{\mathrm{tweedie}}$.
The above analysis shows that the estimate of the mean of the denoising distribution given by the Tweedie’s formula based approach is exact up to error in the score model. Under a particular choice of parameterization and discretization, the probability flow ODE based approach can be made to coincide with the Tweedie’s formula based approach. However, in general the probability flow ODE based approach will introduce additional sources of error in the mean estimate that are not present in the Tweedie’s formula based approach.
@@ ACKNOWLEDGMENTS
The authors would like to thank Dr. Sebastian Reich, Dr. Lars Ruthotto, and Dr. Nick Alger for inspiring discussions on diffusion modeling and randomize-then-optimize algorithms.

### v2

@@ 5.3 STYLIZED PHASE RETRIEVAL STUDY
Table 5 shows the average errors in estimation of the mean and variance, as well as average CMD and MMD errors, for all nine BIPSDA methods tested. Here we first note that, when the score is known analytically, the ‘TC’ variants perform better than the ‘TU’ and ‘ODE’ variants, and on some trials produce strong performance, as evidenced by the provided interdecile ranges. This indicates that the ‘TC’ variants have some promise if a reliable approximation of the second-order score can be estimated using, e.g. the methods in [37]. However, there is wide variance in performance across the trials, and in general the ‘TC’ variants cannot fully overcome the issues observed with the ‘TU’ variants in Figure 2. All of the BIPSDA algorithms we tested thus face performance issues in this problem context.
5.4 COMPUTATIONAL CONSIDERATIONS
Here we examine the computational cost associated with the different BIPSDA algorithms. Table 6 shows the runtimes associated with each tested BIPSDA algorithm across all four studies (learned score regime). For each test-case, we measured the runtime to generate 10,000 samples per trial. The reported runtimes are the average over 100 trials. As can be seen, ‘TU’ variants are consistently faster than their ‘ODE’ counterparts due to the reduced number of score evaluations associated with ‘TU’ variants. However, the dominant computational cost in our problem settings is sampling from the approximate prediction distribution, not approximating the denoising distribution. Here we find that the ‘MAP’ and ‘RTO’ variants are consistently faster than the ‘Lang’ variants. The difference in runtimes is particularly pronounced in the inpainting studies, where the MAP point can be computed efficiently in closed form. Note that the cost of the Langevin dynamics based approaches could be reduced by reducing the number of Langevin iterations, but this may hurt the quality of the generated samples. In general, the computational feasibility of BIPSDA algorithms in a given problem setting is dependent on finding efficient MAP solvers (for the ‘MAP’ and ‘RTO’ variants) or efficient MCMC algorithms (for the ‘Lang’ variants) for the given problem.
Inpainting (Low Noise) Inpainting (High Noise)   X-ray Tomography Phase Retrieval
Lang-TU 26.698(0.879)26.698\;(0.879) 26.938(1.10526.938\;(1.105) 514.829(11.383)514.829\;(11.383) 333.764(20.521)333.764\;(20.521)
Lang-ODE 31.698(0.457)31.698\;(0.457) 31.696(0.259)31.696\;(0.259) 519.479(10.768)519.479\;(10.768) 337.544(19.719)337.544\;(19.719)
MAP-TU $\mathbf{1.217\;(0.011)}$ $\mathbf{1.213\;(0.011)}$ $\mathbf{161.152\;(29.001)}$ $\mathbf{201.837\;(32.167)}$
MAP-ODE 5.590(0.031)5.590\;(0.031) 5.569(0.033)5.569\;(0.033) $\mathbf{174.486\;(56.858)}$ $\mathbf{210.563\;(35.067)}$
RTO-TU 1.381(0.002)1.381\;(0.002) 1.409(0.008)1.409\;(0.008) $\mathbf{171.669\;(23.894)}$ 269.171(26.925)269.171\;(26.925)
RTO-ODE 5.780(0.008)5.780\;(0.008) 5.763(0.027)5.763\;(0.027) $\mathbf{182.396\;(58.243)}$ 273.492(24.678)273.492\;(24.678)
Table 6: Computational time (in seconds) to generate $10,000$ posterior samples (averaged over $100$ trials) with different BIPSDA methods across all four studies (learned score regime). The standard deviation of the runtime is also displayed in parenthesis. All experiments were run on a Nvidia A100 GPU with 80 GB of memory. As can be seen, while computational cost is highly dependent on the specific implementation of each algorithm, in general the ‘MAP’ and ‘RTO’ variants are faster than the ‘Lang’ variants. The speedup is particularly pronounced in the inpainting studies, where the MAP problem can be solved efficiently in closed form.
@@ 6 DISCUSSION AND CONCLUSION
Diffusion annealing based inverse problem solvers are an emerging class of algorithms for solving Bayesian inverse problems using a diffusion model trained on the prior distribution. In this work, we introduced a general framework, Bayesian Inverse Problem Solvers through Diffusion Annealing (BIPSDA), that provides an unified formulation for this class of algorithms. This framework has two key design choices—the denoising distribution approximation and the sampler for the prediction distribution—that can be “mixed and matched” in a flexible manner. The large algorithmic design space of BIPSDA includes the previously introduced DAPS [22] and DiffPIR [23] algorithms, which have both achieved state-of-the-art performance on different image reconstruction problems. Novel approaches can be unveiled through combinations of ideas from DAPS and DiffPIR, as well as the novel techniques proposed here to approximate the denoising distribution and to sample from the prediction distribution. Specifically, the Tweedie Correlated (‘TC’) technique, previously introduced in the context of the hijacking class of diffusion-based Bayesian inverse problem solvers [33], is a novel contribution in the diffusion annealing context. This approach provides a theoretically-sound, hyperparameter free Gaussian approximation to the denoising distribution by leveraging the second moments of the data distribution and the generalized Tweedie formula [37]. The randomize-then-optimize (‘RTO’) technique, which was originally developed as a proposal distribution for Markov chain Monte Carlo methods [24, 26, 25], is also introduced here for the first time to generate fast, approximate samples from the prediction distribution by use of an “off-the-shelf” numerical optimization method.
We also proposed a suite of four benchmark problems (inspired by image inpainting, x-ray tomography, and phase retrieval) for the rigorous assessment of the proposed BIPSDA framework, as well as other diffusion-based posterior samplers. A key feature of these benchmark problems is that they not only have an analytically-known posterior distribution, but also feature an analytical close-form expression for the noisy prior score. This enables analysis of the robustness of algorithms to errors in the score model by examining the idealized scenario of a known-exactly prior score.
@@ 6.1 RESULTS ANALYSIS
The results provide insights into both the impact of specific design choices within our framework and the performance of diffusion-model-based inverse problem solvers generally. Regarding the choice of the denoising distribution approximation, we note that approaches that use the ‘TC’ variant provide the best performance on the two non-linear inverse problems we tested when the prior score is assumed to be known analytically. However, in our implementation this technique is currently only applicable when the score is known exactly. The training of an auxiliary neural network model to approximate higher-order moments of the data distribution is required in the learned score setting [37]. It is also worth noting that in general, the ‘TU’ variant performs better than the ‘ODE’ variant.
Regarding the sampler for the prediction distribution, the results demonstrate that the MAP estimation based approaches, despite lacking firm theoretical foundations, perform well in recovering the global structure of the posterior distribution. However, they systematically underestimate the variance of each of the posterior modes, which is unsurprising given that the ‘MAP’ variants do not actually sample from the prediction distribution. The ‘RTO’ technique, which is equivalent to exact sampling from the prediction distribution when the likelihood function is linear-Gaussian [43], partially resolves this issue and provides better performance than the ‘MAP’ variants on the stylized x-ray tomography problem and inpainting problems. Finally, the ‘Lang’ variants work well when the posterior is unimodal (low noise regime of the stylized inpainting problem and stylized x-ray tomography problem). However, in the case of multimodal posteriors (as in the high noise regime of the stylized inpainting problem), the ‘Lang’ variants struggle to properly incorporate the measurement information. Further, on the phase retrieval problem, Langevin dynamics without Metropolis correction completely fails, and the incorporation of a Metropolis adjustment and preconditioning was required to obtain competitive performance.
@@ 6.2 LIMITATIONS AND FUTURE DIRECTIONS
First, the performance of BIPSDA algorithms, and diffusion-annealing approaches in general, strongly depends on both the choice of algorithm-specific hyperparameters and the accuracy of the underlying pretrained diffusion model. In our work, we carefully controlled for these effects by fixing some of the hyperparameters (e.g. the same sequence of annealing time steps $t_{N_{A}},t_{N_{A}-1},\cdots,t_{0}]$ for the outer loop of Algorithm 1) and by comparing performances using both learned and analytic noisy prior scores. Furthermore, some of the novel algorithms that we propose within the BIPSDA framework achieve similar or superior performance to state-of-the-art methods, such as DAPS, while drastically reducing the number of hyperparameters that the user must provide. Specifically, the ‘TC’ variant, which our work is the first to explore in the contest of diffusion-annealing approaches, provides a hyperparameter-free approximation of the denoising distribution, while on the contrary the ‘ODE’ variant used by DAPS is highly sensitive to the choice of the discretization parameters of the probability flow ODE. Additionally, the ‘RTO’ variant that we proposed here for the first time computes approximate samples from the prediction distribution using “off-the-shelf” deterministic optimization algorithms. Conversely, the ‘Lang’ variant used by DAPS requires careful tuning of the time step size and number of time steps in the Langevin dynamics. Nevertheless, while in this work we followed recommendations in the literature when choosing hyperparameters [22, 54], more work is required to understand the full impact of user provided hyperparameter choices (such as the noise annealing schedule) and how to optimize the accuracy-computational cost trade-off.
Second, the algorithms in our framework were systematically tested on stylized model problems that enable principled performance analysis but have low dimensionality. In the Supplementary Material we also apply algorithms from the BIPSDA family to imaging problems of practical relevance. Proof-of-principle numerical results related to image inpainting (c.f. Fig. S.5 and Table S.1 in the Supplementary Material) demonstrate the computational feasibility of applying algorithms from the BIPSDA family at scale and their ability to produce plausible, diverse samples. However, to systematically evaluate the performance BIPSDA framework in these high-dimensional settings, there remains the need for image reconstruction-relevant, large-scale benchmarks with well-characterized image priors.
Third, it is also of interest to improve the performance of BIPSDA algorithms on challenging non-linear problems like phase retrieval. Here we first note that in this study there were cases where posterior samples numerically diverged due to large error in the learned score in low probability regions of the prior. This could potentially be addressed through the incorporation of second-order score model derivative information into the score training, which has been shown to improve model accuracy in low-density regions in other problem contexts [55]. Further, even with the score known analytically, the BIPSDA algorithms performed poorly on some of the trials, and additional algorithmic innovations are required to address this. In particular, it is of interest to explore modifications to the proposed ‘RTO’ technique, such as apodizing the noise perturbation in early BIPSDA iterations or metropolizing the RTO sampling, to improve performance on challenging problems like this one.
Finally, while the focus of this work was to provide numerical insight on the ability of decoupled noise annealing type approaches to accurately sample from the posterior distribution, the three benchmark problems that we designed are expected to become a useful tool to the community to develop, refine, and rigorously assess existing and novel approaches (including diffusion-type posterior samplers that lie outside of the BIPSDA framework), for solving Bayesian inverse problems with data-driven priors.
@@
7 RELATIONSHIP BETWEEN TWEEDIE FORMULA AND THE PROBABILITY FLOW ODE
In this appendix we analyze the relationship between the Tweedie’s formula and probability flow ODE based approaches for estimation the mean of the denoising distribution $\mathbb{E}_{0\mid t}[\mathbf{m}(0)\mid\mathbf{m}(t)]$.
In the Tweedie’s formula based approach, the estimate of the predicted mean, denoted $\mathbf{m}_{\mathrm{tweedie}}$, is computed using the pretrained score model as
$\mathbf{m}_{\mathrm{tweedie}}=\mathbf{m}(t)+\sigma^{2}(t)s_{\boldsymbol{\theta}^{*}}(\mathbf{m}(t),t),$ (14)
where we have again assumed $\sigma(0)=0$. Since the conditional distribution of $\mathbf{m}(t)$ given $\mathbf{m}(0)$ is Gaussian, Tweedie’s formula is exact and
$\mathbb{E}_{0\mid t}[\mathbf{m}(0)\mid\mathbf{m}(t)]=\mathbf{m}(t)+\sigma^{2}(t)\nabla_{\mathbf{m}(t)}\log\pi_{t}(\mathbf{m}(t)).$
@@ 7 RELATIONSHIP BETWEEN TWEEDIE FORMULA AND THE PROBABILITY FLOW ODE
As can be seen, the ODE and Tweedie’s formula based approaches both update $\mathbf{m}(t)$ in the direction of $s_{\boldsymbol{\theta}^{*}}(\mathbf{m}(t),t)$. However, the magnitude of the update differs by a factor of $d=t\sigma(t)\dot{\sigma}(t)/\sigma^{2}(t)$, which in general will not be equal to one, and so therefore $\mathbf{m}_{\mathrm{ode}}\neq\mathbf{m}_{\mathrm{tweedie}}$. However, under the particular choice of parameterization $\sigma(t)=t$, $d=1$ and the updates coincide.
The above analysis shows that the estimate of the mean of the denoising distribution given by the Tweedie’s formula based approach is exact up to error in the score model. Under a particular choice of parameterization and discretization, the probability flow ODE based approach can be made to coincide with the Tweedie’s formula based approach. However, in general the probability flow ODE based approach will introduce additional approximations in the mean estimate.
@@ ACKNOWLEDGMENTS
The authors would like to thank Drs. Sebastian Reich, Lars Ruthotto, and Nick Alger for inspiring discussions on diffusion modeling and randomize-then-optimize algorithms.